# SWE-Sharp-Bench: A Reproducible Benchmark for C# Software Engineering Tasks

### v1

@@ A-E Extended Results
SWE-agent GPT-4o 26.00 5.11 8.00
Claude Sonnet 3.5 33.60 10.20 19.33
Claude Sonnet 3.7 62.40 14.68 30.67
@@
Claude Sonnet 3.5 53.00 12.73 21.00
Claude Sonnet 3.7 60.60 16.01 28.67
@@ A-E1 Performance by Repository:
OpenHands SWE-Agent Repository GPT-4o GPT-5 Sonnet-3.5 Sonnet-4 GPT-4o GPT-5 Sonnet-3.5 Sonnet-4 dotnet/efcore 0.0 55.6 33.3 44.4 55.6 55.6 44.4 55.6 serilog/serilog 33.3 50.0 33.3 58.3 16.7 41.7 33.3 58.3 spectreconsole/spectre.console 10.0 60.0 40.0 60.0 10.0 60.0 20.0 60.0 SixLabors/ImageSharp 0.0 62.5 25.0 75.0 12.5 62.5 25.0 50.0 autofac/Autofac 16.7 50.0 33.3 50.0 0.0 50.0 33.3 50.0 restsharp/RestSharp 0.0 60.0 20.0 60.0 0.0 60.0 20.0 60.0 dotnet/BenchmarkDotNet 12.5 62.5 25.0 50.0 0.0 62.5 12.5 50.0 devlooped/moq 0.0 50.0 25.0 75.0 0.0 50.0 0.0 75.0 AvaloniaUI/Avalonia 9.8 41.5 24.4 31.7 14.6 36.6 22.0 41.5 StackExchange/StackExchange.Redis 0.0 33.3 0.0 33.3 0.0 33.3 33.3 66.7 App-vNext/Polly 0.0 42.9 21.4 28.6 7.1 42.9 14.3 35.7 MessagePack-CSharp/MessagePack-CSharp 0.0 50.0 5.6 27.8 5.6 44.4 5.6 38.9 ardalis/CleanArchitecture 33.3 33.3 33.3 0.0 0.0 33.3 33.3 0.0 ThreeMammals/Ocelot 0.0 25.0 0.0 50.0 0.0 0.0 0.0 0.0 jellyfin/jellyfin 0.0 33.3 0.0 0.0 0.0 0.0 0.0 33.3 gui-cs/Terminal.Gui 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 JoshClose/CsvHelper 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 OVERALL 8.0 47.3 22.7 40.7 11.3 43.3 20.0 44.7

### v3

@@ A-E Extended Results
SWE-agent GPT-4o 26.00 5.11 11.30
Claude Sonnet 3.5 33.60 10.20 20.00
Claude Sonnet 3.7 62.40 14.68 31.30
@@
Claude Sonnet 3.5 53.00 12.73 22.60
Claude Sonnet 3.7 60.60 16.01 31.30
@@ A-E1 Performance by Repository:
OpenHands SWE-Agent Repository GPT-4o GPT-5 Sonnet-3.5 Sonnet-4 GPT-4o GPT-5 Sonnet-3.5 Sonnet-4 dotnet/efcore 0.0 55.6 33.3 44.4 55.6 55.6 44.4 55.6 serilog/serilog 33.3 50.0 33.3 58.3 16.7 41.7 33.3 58.3 spectreconsole/spectre.console 10.0 60.0 40.0 60.0 10.0 60.0 20.0 60.0 SixLabors/ImageSharp 0.0 62.5 25.0 75.0 12.5 62.5 25.0 50.0 autofac/Autofac 16.7 50.0 33.3 50.0 0.0 50.0 33.3 50.0 restsharp/RestSharp 0.0 60.0 20.0 60.0 0.0 60.0 20.0 60.0 dotnet/BenchmarkDotNet 12.5 62.5 25.0 50.0 0.0 62.5 12.5 50.0 devlooped/moq 0.0 50.0 25.0 75.0 0.0 50.0 0.0 75.0 AvaloniaUI/Avalonia 9.8 41.5 24.4 31.7 14.6 36.6 22.0 41.5 StackExchange/StackExchange.Redis 0.0 33.3 0.0 33.3 0.0 33.3 33.3 66.7 App-vNext/Polly 0.0 42.9 21.4 28.6 7.1 42.9 14.3 35.7 MessagePack-CSharp/MessagePack-CSharp 0.0 50.0 5.6 27.8 5.6 44.4 5.6 38.9 ardalis/CleanArchitecture 33.3 33.3 33.3 0.0 0.0 33.3 33.3 0.0 ThreeMammals/Ocelot 0.0 25.0 0.0 50.0 0.0 0.0 0.0 0.0 jellyfin/jellyfin 0.0 33.3 0.0 0.0 0.0 0.0 0.0 33.3 gui-cs/Terminal.Gui 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 JoshClose/CsvHelper 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 OVERALL 8.0 47.3 22.6 40.6 11.3 43.3 20.0 44.7